OpenStudy (anonymous):

The population size of a country grows exponentially at a rate of 0.95% per year. After how many years will the population have doubled in size?

OpenStudy (anonymous):

@Luigi0210 please help

OpenStudy (anonymous):

Hi luigi ;D

OpenStudy (luigi0210):

hello dwade~ And you don't have any other info given?

OpenStudy (anonymous):

thats literally everything the question is asking

OpenStudy (anonymous):

i don't know why they gave me such little info. im kinda confused

OpenStudy (the_fizicx99):

Well it's below 0.99 so it's a decay... >.>

OpenStudy (luigi0210):

Fizzy can help^

OpenStudy (anonymous):

really? lol but wouldn't that change the question completely?

OpenStudy (the_fizicx99):

o_o

OpenStudy (the_fizicx99):

Yes... why I'm confused ;_;

OpenStudy (anonymous):

copied and pasted that question so nothings a typo

OpenStudy (anonymous):

did i mention my professor is kind of an idiot?

OpenStudy (the_fizicx99):

I honestly don't know, but I haven't done Math in almost 4 days D: So I keep doubting myself more then usual :l @ganeshie8 might know how to solve this :>

OpenStudy (anonymous):

alright cool hopefully he can see his bat signal im in desperate need of assistance right now

OpenStudy (anonymous):

@freckles know whats up with this question?

OpenStudy (luigi0210):

@jim_thompson5910 @whpalmer4 @zepdrix

OpenStudy (anonymous):

luigi think you could answer this one instead? The initial amount of a radioactive substance is 100 grams and is decreasing exponentially at a yearly rate of 10%. In what year will there be half of the substance left?

OpenStudy (anonymous):

@Luigi0210 you there bro?

OpenStudy (luigi0210):

Yea, and I think you just plug that into the decay formula and work backwards http://prntscr.com/3fd8wt So we have \(50=100(.9)^x\) I think. Just solve for x.

OpenStudy (whpalmer4):

No, it's 0.95% growth — that means \((1+\frac{0.95}{100})\) times the previous value each year. It is not a decay.

OpenStudy (the_fizicx99):

Oh I misread that >.<

OpenStudy (whpalmer4):

\[2 = (1+\frac{0.95}{100})^n\]\[\log 2 =n \log 1.0095\]\[n = \frac{\log 2}{\log 1.0095}\]